OpenStudy (anonymous):

Verify 4 csc 2x = 2 csc^2 x tanx Help me please

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

lets take left side first. 4 csc 2x convert csc x to sin x, know how ?

OpenStudy (anonymous):

csc = 1/sin

OpenStudy (anonymous):

how would I write that? The 4 and 2x are confusing me

hartnn (hartnn):

now csc 2x = 1/ sin 2x 4 csc 2x = 4/ sin 2x you know the formula for sin 2x ?

OpenStudy (anonymous):

No

hartnn (hartnn):

sin 2x = 2 sin x cos x plug this in

OpenStudy (anonymous):

4/2 sin x cos x

hartnn (hartnn):

ys, and 4/2 = 2 so for left side we have 2/(sin x cos x) now we will try to simplify the right side, and check whether we get same expression. 2 csc^2 x tanx convert both csc x and tan x in terms of sin x and cos x can u

OpenStudy (anonymous):

we would change csc^2 to cot^2 + 1?

hartnn (hartnn):

since we know csc x = 1/sin x try replacing csc^2 x by 1/ sin^2 x

OpenStudy (anonymous):

2/sin^2x

hartnn (hartnn):

and what about that tan x ?

OpenStudy (anonymous):

sin x/cos x

hartnn (hartnn):

yes, so  what does that simplify to ?